OpenStudy (anonymous):

It takes Kim 8 hours to travel a distance of 360 miles. At this rate, how many miles does Kim travel in 12 hours? 5 minutes... go

OpenStudy (thesecret20111):

Ok try thinking about it like this, if Kim travels 360 miles every 8 hours how many miles would she travel in 1 hour?

OpenStudy (anonymous):

make a proportion... 8/360 = 12/x

OpenStudy (anonymous):

8/360 = 12/x 8x=360*12 x=360*12/8

OpenStudy (anonymous):

It's a contest for us on this site? Very creative way to get others to do your HW for you